# Lightweight posterior construction for gravitational-wave catalogs with the Kolmogorov-Arnold network

### v1

@@ Appendix A GW Events and Sampling Results
Appendix B Population Model and Hyperparameters
In this appendix, we provide detailed description and priors of the hyperparameters used in our population inference. We employ the two-component-primary-mass-ratio and iid-spin models from the GWPopulation package Talbot et al. (2025) for population inference. The two-component-primary-mass-ratio is a power law model for two-dimensional mass distribution. It models the primary mass and the conditional mass ratio as,
@@ Appendix B Population Model and Hyperparameters
Here $\chi_{1}$ and $\chi_{2}$ are the spin magnitudes of the two black holes, and $\theta_{1}$ and $\theta_{2}$ are the spin orientations. $\rm Beta(\cdot)$ denotes the Beta distribution, $\mathcal{N}_{[-1,1]}$ denotes a truncated Gaussian distribution in the range [1,1][-1,1], and $\alpha_{\chi}$ and $\beta_{\chi}$ are the parameters of the Beta distribution. More description and prior choices of the hyperparameters are shown in Table 3.
Table 3: Definitions and prior distributions for the hyperparameters used in our population inference with the GWpopulation package Talbot et al. (2025). The notation $\mathcal{U}(\theta;a,b)$ represents a uniform prior for the parameter $\theta$ over the interval $[a,b]$.

### v2

@@ Appendix B GW Events and Sampling Results
Table 4: Normalized one-dimensional Wasserstein distances for the posteriors of data products of other nine events. Similar to Tab. 2.
Event Type $\mathcal{M}$ $q$ $\chi_{1}$ $\chi_{2}$ $\theta_{1}$ $\theta_{2}$
GW151012 Network Weight 0.012 0.007 0.006 0.008 0.016 0.020
Analytic Expression 0.007 0.011 0.025 0.005 0.016 0.019
GW151226 Network Weight 0.118 0.025 0.010 0.019 0.016 0.036
Analytic Expression 0.116 0.010 0.004 0.004 0.074 0.053
GW170104 Network Weight 0.121 0.014 0.014 0.036 0.021 0.011
Analytic Expression 0.120 0.003 0.005 0.006 0.008 0.069
GW170608 Network Weight 0.066 0.015 0.008 0.014 0.014 0.044
Analytic Expression 0.067 0.018 0.003 0.004 0.020 0.094
GW170729 Network Weight 0.067 0.007 0.007 0.011 0.012 0.009
Analytic Expression 0.022 0.016 0.006 0.007 0.036 0.023
GW170809 Network Weight 0.034 0.018 0.008 0.013 0.012 0.015
Analytic Expression 0.033 0.007 0.008 0.003 0.012 0.051
GW170814 Network Weight 0.031 0.007 0.004 0.015 0.010 0.026
Analytic Expression 0.030 0.005 0.007 0.010 0.008 0.047
GW170818 Network Weight 0.024 0.007 0.009 0.009 0.012 0.024
Analytic Expression 0.022 0.011 0.005 0.006 0.008 0.046
GW170823 Network Weight 0.074 0.012 0.004 0.024 0.023 0.012
Analytic Expression 0.076 0.009 0.005 0.010 0.014 0.010
Appendix C A Full 15-dimensional Posterior Reconstruction for GW150914
In this appendix, we present the reconstruction and compression results for the full 15-dimensional posterior distribution of GW150914 using the KAN-based density estimator. Apart from the input dimension, the network architecture is essentially the same as in the 6-dimensional case, namely a compact two-layer network. Each conditional density is modeled with 5 Gaussian components, and each unmasked edge contains 23 cubic B-spline basis functions. Despite the increased dimensionality, the trained network remains compact. The stored network weights occupy only about 0.5 megabytes, compared with the 16.9 megabytes required to store the original posterior samples, corresponding to a $34\times$ compression ratio. The resulting reconstructed posterior distribution is shown in Fig. 12, which exhibits close agreement with the original distribution, demonstrating the effectiveness and scalability of our KAN-based density estimator.
Figure 12: Marginalized one- and two-dimensional distributions, comparing samples generated with the neural network weights (orange) and raw samples (blue) for the full 15-dimensional posterior of GW150914. Contour lines in the two-dimensional joint distributions delineate the $50\%$ and $90\%$ credible regions. In this figure, we have used $\Delta t_{c}\equiv t_{c}-1126259462$ s.
Appendix D Population Model and Hyperparameters
In this appendix, we provide detailed description and priors of the hyperparameters used in our population inference. We employ the two-component-primary-mass-ratio and iid-spin models from the GWPopulation package Talbot:2024yqw for population inference. The two-component-primary-mass-ratio is a power law model for two-dimensional mass distribution. It models the primary mass and the conditional mass ratio as,
@@ Appendix D Population Model and Hyperparameters
Here $\chi_{1}$ and $\chi_{2}$ are the spin magnitudes of the two black holes, and $\theta_{1}$ and $\theta_{2}$ are the spin orientations. $\rm Beta(\cdot)$ denotes the Beta distribution, $\mathcal{N}_{[-1,1]}$ denotes a truncated Gaussian distribution in the range [1,1][-1,1], and $\alpha_{\chi}$ and $\beta_{\chi}$ are the parameters of the Beta distribution. More description and prior choices of the hyperparameters are shown in Tab. 5. We adopt the prior ranges following the GWPopulation package [^3] https://colmtalbot.github.io/gwpopulation/examples as a test case to validate the effectiveness of our method. More diverse prior choices can be found in previous studies LIGOScientific:2020kqk; KAGRA:2021duu.
Table 5: Definitions and prior distributions for the hyperparameters used in our population inference with the GWpopulation package Talbot:2024yqw. The notation $\mathcal{U}(\theta;a,b)$ represents a uniform prior for the parameter $\theta$ over the interval $[a,b]$.